OpenStudy (anonymous):

@amberk123 @myininaya

OpenStudy (anonymous):

@cookiimonster27 Instead of re-drawing it you can also right click and save the image then click the attach file button.

OpenStudy (anonymous):

is it the same question hun

OpenStudy (anonymous):



myininaya (myininaya):

that one is on the outside of the sqrt right?

OpenStudy (anonymous):

Whats the solution

OpenStudy (anonymous):

\[\sqrt{3x-5}=\sqrt{4x-3-1}\]

OpenStudy (anonymous):

I Think that you have to do the same to each side.

OpenStudy (anonymous):

\[\sqrt{3x-5}=\sqrt{4x-4}\]

OpenStudy (anonymous):

yes thats it @amberk123

myininaya (myininaya):

\[\sqrt{3x-5}=\sqrt{4x-4}\] see what happens if you square both sides

OpenStudy (aum):

Are you sure the minus 1 on the right hand side is not outside the radical sign?

OpenStudy (anonymous):

@aum yes im sure

OpenStudy (anonymous):

oh wait your its its outside @aum

OpenStudy (aum):

Aha! It is too simple if it is inside!

myininaya (myininaya):

so then we have this: \[\sqrt{3x-5}=\sqrt{4x-3}-1 \] ?

OpenStudy (anonymous):

yes

myininaya (myininaya):

still need to square both sides

myininaya (myininaya):

what do you get when you square both sides?

OpenStudy (anonymous):

how do i do it ?

myininaya (myininaya):

\[(\sqrt{3x-5})^2=(\sqrt{4x-3}-1)^2 \] this gives you?

myininaya (myininaya):

for the right hand side please do not say (a-b)^2=a^2-b^2 because it doesn't (a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (anonymous):

sqrt(3 x-5)-sqrt(4 x-3) = -1

OpenStudy (anonymous):

how ? @amberk123

OpenStudy (anonymous):

okay i will show u

OpenStudy (anonymous):

\[\sqrt{3x-5+1}=\sqrt{4x-3}\] \[\sqrt{3x-5}-\sqrt{4x-3=-1}\]

myininaya (myininaya):

Lets have cookie work for the solution a little.

myininaya (myininaya):

first of all do you know what this equals cookie \[(\sqrt{9})^2=\]?

myininaya (myininaya):

the reason I'm asking is because this is going to go right back to how to answer \[(\sqrt{3x-5})^2\]

OpenStudy (anonymous):

she offline